OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (- 3, - 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 6x -

OpenStudy (campbell_st):

well using the zeros the polynomial is and a is a constant y = a(x +8)(x - 2) so y = a(x^2 + 6x - 16) you now need to find the value of a by substituting the vertex x= -3, y = -25 \[-25 = a((-3)^2 + 6(-3) - 16) \] -25 = a(9 - 18 - 16) a = 1 so the equation is \[y = x^2 + 6x - 16\]